OpenStudy (coledrf):

i need help please!!!!! simplify into one fraction 8x/x-8 - 2/x-8

geerky42 (geerky42):

Both fractions already have same denominator

geerky42 (geerky42):

\(\dfrac{a}{c} + \dfrac{b}{c} \rightarrow \dfrac{a+b}{c}\)

OpenStudy (anonymous):

\(\Large \frac{8}{x-8}-\frac{2}{x-8}=\frac{8-2}{x-8}=\frac{6}{x-8}\)